A student obtained a score of 45 in Math and a score of 60 in Science. If the mean and standard deviation of the math scores are 26.4 and 7.2, respectively, while that of science are 42.5 and 12.6,


1
Expert's answer
2022-04-19T04:40:05-0400

Math: "z_m=\frac{45-26.4}{7.2}=2.58."

Science: "z_s=\frac{60-42.5}{12.6}=1.39."

Since the z-score for Math is higher, the student is more successful in this subject.